# Towards Automated Identification of Violation Symptoms of Architecture Erosion

### v3

@@ 4.1 RQ1: Identifying Violation Symptoms
TABLE II: Performance comparison among ML-based classifiers
@@
TABLE III: Performance comparison among DL-based classifiers
@@
TABLE IV: Performance comparison of the classifiers based on the fastText model with different dimension size
@@
TABLE V: Performance comparison of ensemble classifiers
@@
TABLE VI: Performance comparison among ensemble ML-based classifiers with three word embeddings
@@
For ML-based classifiers, as shown in Table II, we trained the five types of ML classifiers with three word embedding techniques (i.e., word2vec, fastText, and GloVe), leading to 15 classifiers (i.e., the five ML models times three word embeddings). We calculated the average values of the classifiers over the three pre-trained word embeddings. For each classifier based on three word embeddings, we underlined the best result of each metric (horizontal comparison); for each metric of the classifiers (precision, recall, F1, and accuracy), we marked the best metric result of each classifier on three word embeddings in bold (vertical comparison). From the results, we can see that the kNN classifiers have better scores of Precision than other classifiers with an average Precision of 0.816. As for the remaining metrics, the SVM classifiers achieve relatively better performance on nearly all metrics than other classifiers, with an average Recall of 0.801, F1-score of 0.779, and Accuracy of 0.773. Specifically, the SVM classifier based on word2vec performed the best with a Precision of 0.789, a Recall of 0.828, an F1-score of 0.808, and an Accuracy of 0.803.
For DL-based classifiers, as shown in Table III, we trained one type of classifier, i.e., TextCNN, based on the vocabulary set from our dataset and three word embedding techniques (i.e. in total four classifiers) as mentioned in Section 3.3.1. We can see that TextCNN_FT performs the best compared with the other three DL models.
@@
From the results in Tables II and III, we can also observe the differences between the classifiers based on three pre-trained word embedding models. As mentioned before, for each classifier based on three word embeddings, we underlined the best result of each classifier on each metric (horizontal comparison). When using the word2vec pre-trained word embedding, most of the ML classifiers have relatively better performance on Precision and Accuracy. Most of the ML classifiers with the fastText model have fairly better F1-scores. Besides, we found that the TextCNN classifier based on the fastText pre-trained word embedding achieves the best performance on all metrics except for Recall, with a Precision of 0.728, a Recall of 0.549, an F1-score of 0.667, and an Accuracy of 0.672.
In addition, we can see in Table III that the overall performance of the three TextCNN classifiers with pre-trained word embeddings outperforms the TextCNN classifier based on the vocabulary set from our dataset. This is because pre-trained word embedding models contain more features due to the large textual data used for generating word vectors, and they have been proven to be invaluable for improving the performance in NLP tasks [55, 56].
In general, the three pre-trained word embedding models have an impact on the performance of the DL-based classifiers. As shown in Table III, when we compared their classification performance, in most cases, the classifiers with the fastText model can achieve relatively good performance.
@@
According to the results of RQ1.2, the fastText model can help to generate better results than the other two pre-trained word embedding models. Therefore, to further explore the impact of dimension sizes of the pre-trained word embedding models, we trained the classifiers based on the fastText model and chose three common dimension sizes: 100, 200, and 300. In Table IV, similarly as before, we underlined the best result of each classifier on each metric (horizontal comparison). We can see that five of the six ML/DL classifiers can generate better F1-scores based on the fastText model with 200-dimensional word embedding, compared with 100 and 300-dimensional models.
@@
4.2 RQ2: Improving Performance with Voting
@@
To answer RQ2, we employed an ensemble learning technique, i.e., majority voting strategy, to build an ensemble classifier. We compared the performance of the individual classifiers with the ensemble classifiers. More specifically, we explored to what extent combining one type of classifier based on three word embedding techniques can achieve better performance, as shown in Table V and Table VI.
@@
In Table V, the “Mean” column presents the Average values of each metric of each classifier on three word embeddings from Table II and Table III. The “Best” column denotes the corresponding classifier with the best performance from Table II and Table III. The “Voting” column refers to the performance of the ensemble classifier (based on classifiers with three word embeddings) on each metric after utilizing a majority voting strategy. For example, the Voting values of the SVM classifiers in Table V refer to the voting results among the three individual classifiers SVM_w2v, SVM_FT, and SVM_GloVe. “Impro_M” and “Impro_B” represent the improved results of the corresponding ensemble classifiers compared with the results in “Mean” and “Best” columns, respectively.
@@
According to Table V, when we utilized the voting strategy for each ML classifier based on the three word embeddings (i.e., word2vec, fastText, GloVe), we can see that all ensemble classifiers perform better than the mean values of individual classifiers, with F1-score improvement from 0.48% to 13.10%. Even compared with the best-performing classifiers, most of the ensemble classifiers outperform the individual ML/DL classifiers, except for kNN. Moreover, the ensemble TextCNN of the three TextCNN classifiers based on pre-trained word embeddings (see Table III) does not perform as well as TextCNN_FT (i.e., the best one) on Precision, F1-score, and Accuracy. This means that TextCNN_FT with the fastText pre-trained word embedding performs best compared with the other three TextCNN classifiers, and the voting strategy does not have an obvious improvement for the trained DL classifiers.
@@
In addition, as shown in Table VI, we utilized the voting strategy for the five ML classifiers based on each of the three word embeddings. For example, ML_word2vec is the ensemble classifier of the five ML classifiers based on the word2vec word embedding (i.e., SVM_w2v, LR_w2v, NB_w2v, DT_w2v, and kNN_w2v). Similarly as before, the “Mean” column refers to the average metric results of ML classifiers (see the last row in Table II); the “Best” column presents the best-performing classifiers in Table II; the “Voting” column denotes the performance results of the ensemble classifiers after voting. It is clear that the three ensemble classifiers outperform the average performance results of the individual classifiers, with improvements from 3.11% to 7.98%. Even compared with the best-performing individual classifiers (i.e., the maximum values of the voting samples), the ensemble classifier ML_fastText can still demonstrate performance improvement on all metrics.
@@ 4.3 RQ3: Validation in Practice
Overall, we planned to send customized survey invitation emails to 200 code reviewers who were involved in the discussion of the collected review comments concerning architecture violations. However, we could only derive valid email addresses for 169 of them. We finally obtained 24 valid responses to our survey and interviewed 6 participants after sending 169 survey invitation emails and reminder emails; the survey response rate of 14.2%, is a reasonable response rate considering that the general response rate is around 5% in empirical software engineering research [57]. Table VII shows the demographic information of the 24 respondents. Twenty two of the respondents have more than 10 years of experience in software development, while two of them have 6-10 years of experience. Among the respondents, 50.0% (12) are software engineers, 33.3% (8) are developers, and 16.7% (4) are architects. Note that, software engineers usually have a broader work range than developers; for example, the former may work on full-stack development while the latter could deal only with the back end. Regarding the size of their companies, ten of the respondents reported that their companies have more than 1000 employees, five respondents worked in companies with employees between 501 and 1000, six respondents in companies with employees between 101 and 500, and three respondents in companies with employees between 21 and 100. Besides, the respondents worked in a variety of application domains.
@@
TABLE VII: Demographic information of the survey participants and interviewees
@@
Figure 4: Practitioners’ responses to statements regarding the usefulness of the trained models
5 Discussion
5.1 Interpretation of Results
5.1.1 RQ1: Identifying Violation Symptoms
Our trained ML/DL classifiers can identify violation symptoms in code review comments. The results of RQ1.1 indicate that certain factors such as the type of classifier, the pre-trained word embeddings, and hyperparameters influence the performance of classifiers. Specifically, we found that the SVM classifier with word2vec outperforms other ML and DL-based classifiers. We also found that DL classifiers do not always have better performance than ML classifiers; we note that training DL classifiers takes much longer time than training ML classifiers, as DL classifiers have much more complex network architecture and more parameters. In recent years, some studies that leveraged DL-based techniques for software engineering tasks have achieved better performance that all other state-of-the-art techniques (e.g., ML-based techniques) [44]. However, the achieved results imply that ML classifiers might be enough for certain text classification tasks with small-size datasets, especially as the trained DL classifiers do not show performance improvement in such a scenario. This finding is also observed in previous studies [58, 59]: for certain software engineering tasks, simple ML-based approaches are capable of achieving equal or even better performance than DL-based approaches within less time. Thus, we advise researchers to not blindly apply DL-based techniques and consider simple approaches first.
@@ 5.1.1 RQ1: Identifying Violation Symptoms
From the results of RQ1.2, we can see that the TextCNN classifier based on the fastText pre-trained word embedding can perform better than other DL models. We thus emphasize that the fastText pre-trained technique has a better capacity for training well-performed DL models; this also aligns with the finding in the work of Sesari et al. [60], that is, the fastText pre-trained word embedding model has less bias compared to word2vec and GloVe.
Besides, the results of RQ1.3 show that 200-dimensional pre-trained word embedding models outperform 100 and 300-dimensional models. Our results are slightly different from previous studies (e.g., [61, 62]) where the 300-dimensional word embedding model has better performance. One possible reason is that the previous studies only measured the average results of classifiers with different dimensional word embedding models, and the average results might be influenced by the individual results. Our results present and compare the classification performance of each ML/DL classifier, which could be more accurate.
@@
5.1.2 RQ2: Improving Performance with Voting
Evaluating the performance of an ensemble classifier offers insight into the performance improvement compared with individual classifiers. For voting-based approaches, an accurate output is attained only when a majority of the constituent models generate the correct output [63]. Our results provide further empirical evidence for the efficacy of the majority voting strategy regarding improving the performance of text classification. This technique effectively combines the predictions of multiple classifiers to arrive at a final decision, resulting in improved overall classification performance. Nevertheless, we also observed that such performance improvement is not obvious in TextCNN in our case. One possible reason is that we only get an ensemble DL classifier based on three individual TextCNN models, and that causes the voting results to be particularly susceptible to the influence of single results. We argue that the voting strategy remains valuable to minimize potential prediction errors if there are more outputs for voting.
@@ 5.1.3 RQ3: Validation in Practice
Finding 4: Although there are concerns about the correctness of the identified architecture violations, the positive responses from participants suggest that practitioners are receptive to the use of automated techniques for identifying architecture violations during code review.
5.2 Implications
@@
6 Threats to Validity
In this section, we discuss the potential threats to the validity related to our study and the adopted measures mitigating these threats by following the guidelines proposed by Wohlin et al. [64].
6.1 Construct validity
Construct validity concerns the connection between operational measures and the studied subjects. The first potential threat to construct validity is the suitability of our selected ML/DL algorithms, feature extraction techniques, and metrics. As mentioned by Peter et al. [65], it is not possible to use an exhaustive approach to test all learning models. We selected five common ML algorithms, one DL algorithm, and three popular pre-trained word embedding models, as they are all widely used to conduct text classification. We admit that using different algorithms and pre-trained word embedding models may generate varying results, despite their popularity. We cannot completely mitigate this threat. In addition, the four metrics (recall, precision, F1-score, and accuracy) are widely used to evaluate the performance of various automatic software engineering techniques (see Section 7.3). Therefore, the metrics pose little threat to the construct validity.
@@ 6.1 Construct validity
6.2 External validity
External validity pertains to the generalizability of our experiment results and findings. We trained the ML and DL-based classifiers with our collected dataset regarding architecture violations. Selecting different datasets and experimental settings might influence the generalizability of (part of) our findings. We acknowledge that the chosen open-source projects are not representative of all projects from the OpenStack and Qt platforms, and our findings may not be generalized to diverse projects, particularly those from different domains such as closed-source projects or projects on other platforms like Apache. Furthermore, the participants of the survey and interviews are contributors to the selected projects, therefore, the findings of RQ3 might not be generalized to other projects. To support further generalizability, we have made our dataset [36] available to enable other researchers and practitioners to replicate, reuse, and extend our study.
6.3 Reliability
Reliability reflects the replicability of a study regarding yielding the same or similar results when other researchers reproduce this study. One potential threat to reliability comes from the experimental settings (e.g., input data and hyperparameters for training classifiers) used in this work. We used the dataset from our previous study [11] that had publicly shared the labeled data and corresponding scripts for collecting data. Besides, we also shared the source code in our replication package [36] to facilitate reproduction and validation. Another potential threat might be from the survey and interview design. To mitigate the threat, we provided the templates of our survey, semi-structured interviews, and customized emails in our replication package [36] to encourage further research in this area.
7 Related Work

### v4

@@ 4.1. RQ1: Identifying Violation Symptoms
Table 2. Performance comparison among ML-based classifiers
@@
Table 3. Performance comparison among DL-based classifiers
@@
Table 4. Performance comparison of the classifiers based on the fastText model with different dimension size
@@
Table 5. Performance comparison of ensemble classifiers
@@
Table 6. Performance comparison among ensemble ML-based classifiers with three word embeddings
@@
For ML-based classifiers, as shown in Table 2, we trained the five types of ML classifiers with three word embedding techniques (i.e., word2vec, fastText, and GloVe), leading to 15 classifiers (i.e., the five ML models times three word embeddings). We calculated the average values of the classifiers over the three pre-trained word embeddings. For each classifier based on three word embeddings, we underlined the best result of each metric (horizontal comparison); for each metric of the classifiers (precision, recall, F1, and accuracy), we marked the best metric result of each classifier on three word embeddings in bold (vertical comparison). From the results, we can see that the kNN classifiers have better scores of Precision than other classifiers with an average Precision of 0.816. As for the remaining metrics, the SVM classifiers achieve relatively better performance on nearly all metrics than other classifiers, with an average Recall of 0.801, F1-score of 0.779, and Accuracy of 0.773. Specifically, the SVM classifier based on word2vec performed the best with a Precision of 0.789, a Recall of 0.828, an F1-score of 0.808, and an Accuracy of 0.803.
For DL-based classifiers, as shown in Table 3, we trained one type of classifier, i.e., TextCNN, based on the vocabulary set from our dataset and three word embedding techniques (i.e. in total four classifiers) as mentioned in Section 3.3.1. We can see that TextCNN_FT performs the best compared with the other three DL models.
@@
From the results in Tables 2 and 3, we can also observe the differences between the classifiers based on three pre-trained word embedding models. As mentioned before, for each classifier based on three word embeddings, we underlined the best result of each classifier on each metric (horizontal comparison). When using the word2vec pre-trained word embedding, most of the ML classifiers have relatively better performance on Precision and Accuracy. Most of the ML classifiers with the fastText model have fairly better F1-scores. Besides, we found that the TextCNN classifier based on the fastText pre-trained word embedding achieves the best performance on all metrics except for Recall, with a Precision of 0.728, a Recall of 0.549, an F1-score of 0.667, and an Accuracy of 0.672.
In addition, we can see in Table 3 that the overall performance of the three TextCNN classifiers with pre-trained word embeddings outperforms the TextCNN classifier based on the vocabulary set from our dataset. This is because pre-trained word embedding models contain more features due to the large textual data used for generating word vectors, and they have been proven to be invaluable for improving the performance in NLP tasks (Qi2018pwe, ; Radford2018ilu, ).
In general, the three pre-trained word embedding models have an impact on the performance of the DL-based classifiers. As shown in Table 3, when we compared their classification performance, in most cases, the classifiers with the fastText model can achieve relatively good performance.
@@
According to the results of RQ1.2, the fastText model can help to generate better results than the other two pre-trained word embedding models. Therefore, to further explore the impact of dimension sizes of the pre-trained word embedding models, we trained the classifiers based on the fastText model and chose three common dimension sizes: 100, 200, and 300. In Table 4, similarly as before, we underlined the best result of each classifier on each metric (horizontal comparison). We can see that five of the six ML/DL classifiers can generate better F1-scores based on the fastText model with 200-dimensional word embedding, compared with 100 and 300-dimensional models.
@@
4.2. RQ2: Improving Performance with Voting
@@
To answer RQ2, we employed an ensemble learning technique, i.e., majority voting strategy, to build an ensemble classifier. We compared the performance of the individual classifiers with the ensemble classifiers. More specifically, we explored to what extent combining one type of classifier based on three word embedding techniques can achieve better performance, as shown in Table 6 and Table 6.
@@
In Table 6, the “Mean” column presents the Average values of each metric of each classifier on three word embeddings from Table 2 and Table 3. The “Best” column denotes the corresponding classifier with the best performance from Table 2 and Table 3. The “Voting” column refers to the performance of the ensemble classifier (based on classifiers with three word embeddings) on each metric after utilizing a majority voting strategy. For example, the Voting values of the SVM classifiers in Table 6 refer to the voting results among the three individual classifiers SVM_w2v, SVM_FT, and SVM_GloVe. “Impro_M” and “Impro_B” represent the improved results of the corresponding ensemble classifiers compared with the results in “Mean” and “Best” columns, respectively.
@@
According to Table 6, when we utilized the voting strategy for each ML classifier based on the three word embeddings (i.e., word2vec, fastText, GloVe), we can see that all ensemble classifiers perform better than the mean values of individual classifiers, with F1-score improvement from 0.48% to 13.10%. Even compared with the best-performing classifiers, most of the ensemble classifiers outperform the individual ML/DL classifiers, except for kNN. Moreover, the ensemble TextCNN of the three TextCNN classifiers based on pre-trained word embeddings (see Table 3) does not perform as well as TextCNN_FT (i.e., the best one) on Precision, F1-score, and Accuracy. This means that TextCNN_FT with the fastText pre-trained word embedding performs best compared with the other three TextCNN classifiers, and the voting strategy does not have an obvious improvement for the trained DL classifiers.
@@
In addition, as shown in Table 6, we utilized the voting strategy for the five ML classifiers based on each of the three word embeddings. For example, ML_word2vec is the ensemble classifier of the five ML classifiers based on the word2vec word embedding (i.e., SVM_w2v, LR_w2v, NB_w2v, DT_w2v, and kNN_w2v). Similarly as before, the “Mean” column refers to the average metric results of ML classifiers (see the last row in Table 2); the “Best” column presents the best-performing classifiers in Table 2; the “Voting” column denotes the performance results of the ensemble classifiers after voting. It is clear that the three ensemble classifiers outperform the average performance results of the individual classifiers, with improvements from 3.11% to 7.98%. Even compared with the best-performing individual classifiers (i.e., the maximum values of the voting samples), the ensemble classifier ML_fastText can still demonstrate performance improvement on all metrics.
@@ 4.3. RQ3: Validation in Practice
Overall, we planned to send customized survey invitation emails to 200 code reviewers who were involved in the discussion of the collected review comments concerning architecture violations. However, we could only derive valid email addresses for 169 of them. We finally obtained 24 valid responses to our survey and interviewed 6 participants after sending 169 survey invitation emails and reminder emails; the survey response rate of 14.2%, is a reasonable response rate considering that the general response rate is around 5% in empirical software engineering research (gaese2008CH1, ). Table 7 shows the demographic information of the 24 respondents. Twenty two of the respondents have more than 10 years of experience in software development, while two of them have 6-10 years of experience. Among the respondents, 50.0% (12) are software engineers, 33.3% (8) are developers, and 16.7% (4) are architects. Note that, software engineers usually have a broader work range than developers; for example, the former may work on full-stack development while the latter could deal only with the back end. Regarding the size of their companies, ten of the respondents reported that their companies have more than 1000 employees, five respondents worked in companies with employees between 501 and 1000, six respondents in companies with employees between 101 and 500, and three respondents in companies with employees between 21 and 100. Besides, the respondents worked in a variety of application domains.
@@
Table 7. Demographic information of the survey participants and interviewees
@@
Figure 4. Practitioners’ responses to statements regarding the usefulness of the trained models
5. Discussion
5.1. Interpretation of Results
5.1.1. RQ1: Identifying Violation Symptoms
Our trained ML/DL classifiers can identify violation symptoms in code review comments. The results of RQ1.1 indicate that certain factors such as the type of classifier, the pre-trained word embeddings, and hyperparameters influence the performance of classifiers. Specifically, we found that the SVM classifier with word2vec outperforms other ML and DL-based classifiers. We also found that DL classifiers do not always have better performance than ML classifiers; we note that training DL classifiers takes much longer time than training ML classifiers, as DL classifiers have much more complex network architecture and more parameters. In recent years, some studies that leveraged DL-based techniques for software engineering tasks have achieved better performance that all other state-of-the-art techniques (e.g., ML-based techniques) (Yang2022pmse, ). However, the achieved results imply that ML classifiers might be enough for certain text classification tasks with small-size datasets, especially as the trained DL classifiers do not show performance improvement in such a scenario. This finding is also observed in previous studies (Liu2018nmt, ; Fu2017eoh, ): for certain software engineering tasks, simple ML-based approaches are capable of achieving equal or even better performance than DL-based approaches within less time. Thus, we advise researchers to not blindly apply DL-based techniques and consider simple approaches first.
@@ 5.1.1. RQ1: Identifying Violation Symptoms
From the results of RQ1.2, we can see that the TextCNN classifier based on the fastText pre-trained word embedding can perform better than other DL models. We thus emphasize that the fastText pre-trained technique has a better capacity for training well-performed DL models; this also aligns with the finding in the work of Sesari et al. (Sesari2022pte, ), that is, the fastText pre-trained word embedding model has less bias compared to word2vec and GloVe.
Besides, the results of RQ1.3 show that 200-dimensional pre-trained word embedding models outperform 100 and 300-dimensional models. Our results are slightly different from previous studies (e.g., (Ren2019nndSATD, ; Li2022SATD, )) where the 300-dimensional word embedding model has better performance. One possible reason is that the previous studies only measured the average results of classifiers with different dimensional word embedding models, and the average results might be influenced by the individual results. Our results present and compare the classification performance of each ML/DL classifier, which could be more accurate.
@@
5.1.2. RQ2: Improving Performance with Voting
Evaluating the performance of an ensemble classifier offers insight into the performance improvement compared with individual classifiers. For voting-based approaches, an accurate output is attained only when a majority of the constituent models generate the correct output (hansen1990nne, ). Our results provide further empirical evidence for the efficacy of the majority voting strategy regarding improving the performance of text classification. This technique effectively combines the predictions of multiple classifiers to arrive at a final decision, resulting in improved overall classification performance. Nevertheless, we also observed that such performance improvement is not obvious in TextCNN in our case. One possible reason is that we only get an ensemble DL classifier based on three individual TextCNN models, and that causes the voting results to be particularly susceptible to the influence of single results. We argue that the voting strategy remains valuable to minimize potential prediction errors if there are more outputs for voting.
@@ 5.1.3. RQ3: Validation in Practice
Finding 4: Despite a few concerns about the accuracy of the identified architecture violations, the majority of positive feedback from participants indicates that practitioners are receptive to the use of automated techniques for identifying architecture violations during code review.
5.2. Implications
@@
6. Threats to Validity
In this section, we discuss the potential threats to the validity related to our study and the adopted measures mitigating these threats by following the guidelines proposed by Wohlin et al. (Wohlin2012ESE, ).
6.1. Construct validity
Construct validity concerns the connection between operational measures and the studied subjects. The first potential threat to construct validity is the suitability of our selected ML/DL algorithms, feature extraction techniques, and metrics. As mentioned by Peter et al. (peters2017text, ), it is not possible to use an exhaustive approach to test all learning models. We selected five common ML algorithms, one DL algorithm, and three popular pre-trained word embedding models, as they are all widely used to conduct text classification. We admit that using different algorithms and pre-trained word embedding models may generate varying results, despite their popularity. We cannot completely mitigate this threat. In addition, the four metrics (recall, precision, F1-score, and accuracy) are widely used to evaluate the performance of various automatic software engineering techniques (see Section 7.3). Therefore, the metrics pose little threat to the construct validity.
@@ 6.1. Construct validity
6.2. External validity
External validity pertains to the generalizability of our experiment results and findings. We trained the ML and DL-based classifiers with our collected dataset regarding architecture violations. Selecting different datasets and experimental settings might influence the generalizability of (part of) our findings. We acknowledge that the chosen open-source projects are not representative of all projects from the OpenStack and Qt platforms, and our findings may not be generalized to diverse projects, particularly those from different domains such as closed-source projects or projects on other platforms like Apache. Furthermore, the participants of the survey and interviews are contributors to the selected projects, therefore, the findings of RQ3 might not be generalized to other projects. To support further generalizability, we have made our dataset (onlinepackage_TSE, ) available to enable other researchers and practitioners to replicate, reuse, and extend our study.
6.3. Reliability
Reliability reflects the replicability of a study regarding yielding the same or similar results when other researchers reproduce this study. One potential threat to reliability comes from the experimental settings (e.g., input data and hyperparameters for training classifiers) used in this work. We used the dataset from our previous study (Li2023wvs, ) that had publicly shared the labeled data and corresponding scripts for collecting data. Besides, we also shared the source code in our replication package (onlinepackage_TSE, ) to facilitate reproduction and validation. Another potential threat might be from the survey and interview design. To mitigate the threat, we provided the templates of our survey, semi-structured interviews, and customized emails in our replication package (onlinepackage_TSE, ) to encourage further research in this area.
7. Related Work